solve system linear equations by addition method
4x-3y=1
8x+5y=13
please solve and show work...thanks

Found 2 solutions by jorel1380, oberobic:
Answer by jorel1380(3719)

4x-3y=1
8x+5y=13
8x-6y=2
11y=11
y=1
x=1..

Answer by oberobic(2304)

4x -3y = 1
8x + 5y = 13
.
Multiply the first equation by -2, then add.
-8x +6y = -2
8x +5y = 13
11y = 11
y = 1
.
Substitute into the first equation to find x.
.
4x -3 = 1
4x = 4
x = 1
.
Check by substituting into the second equation
8(1) +5(1) = 13
True.
.
Answer: x=1, y=1
.
Done.